2. (15 points) Prove that for a positive integer n, the number gcd (n + 1, na — n + 1) is equal either to 1 or to 3.
Answer #1

ged(n+1, ²-n+1) = Them alm+1), a 1 (2-n+1) and al MA132 dl (n= n+1) >> dl{n+1)*--(-41)} => dl (n = 2n +4-mano-1) 7 dl 3m d/ (geden, n+1) = 1 an I integer uv such that mu 6V=1 Again d 1(n+1) = (n + 1) = ap for some integer P. There fore from a mut dev